[Brahms-l] Shift report 20050310 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Mar 10 2005 - 23:57:43 EST
Shifter leader was kh
Shifters were 


Had the first 63 GeV physics run.  It was "declared" at about 17:45

After we had beam, Flemming made some slight adjustments to the RHIC Clock to have it overlap with the triggers.

Started running at 18:30.  Runs achieved:
Run      tr3     tr6
13674   58k     149k
13675   44k     120k
13676   35k      99k
13677    6k      19k
13678   25k      72k

No real detector problems.  One operator problem (me!)  Runs 13674 through 13677 were not written to HPSS (oversight on my part)  They will need to be transferred into HPSS manually.  A procedure exists for that although I hear it is tedious.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 10 23:58:06 2005

This archive was generated by hypermail 2.1.8 : Thu Mar 10 2005 - 23:58:21 EST